### v1

@@ 7 Conclusion
This paper shows how a popular way of assessing causal mechanisms—the intermediate outcome test—relies on strong and often undisclosed assumptions. In these tests, authors present the average effects of treatment on the mediator and often argue that these effects show plausible causal pathways of the overall effect of the treatment on the outcome. However, our analysis shows that IOTs can neither establish nor rule out indirect effects through a mediator under randomization of treatment alone. We showed that IOTs can rule out indirect effects if the effect of treatment on the mediator can only move units in one direction, which is a very strong monotonicity assumption. Furthermore, IOTs cannot establish an indirect effect even if we assume a positive average effect of the mediator on the outcome. We derive sharp nonparametric bounds for the indirect effect under these various assumptions and estimate these bounds in two empirical applications. Finally, we also discussed how causal mechanisms might be assessed in situations where the assumptions needed to make inferences about indirect effects are questionable.
We believe there is much future research to be completed in the field of causal mechanisms. In particular, future research should attempt to unify different accounts of “mechanisms” in a single broad framework. Currently, mechanisms are sometimes used interchangeably with “indirect effects,” but other times, scholars use mechanisms to refer to a theoretical construct that can explain an effect without reference to any mediation at all. In epidemiology, there is a tradition of defining causal mechanisms in the “sufficient cause framework,” where authors seek to find sets of variable levels that are sufficient to produce a particular outcome (VanderWeele, 2009). Connecting the ideas of mediation and mechanism to the theoretical underpinnings of the social sciences would help scholars develop better empirical tests of their theories.

### v2

@@ 7 Conclusion
This paper shows how a popular way of assessing causal mechanisms—the intermediate outcome test—relies on strong and often undisclosed assumptions. In these tests, authors present the average effects of treatment on the mediator and often argue that these effects show plausible causal pathways of the overall effect of the treatment on the outcome. However, our analysis shows that IOTs can neither establish nor rule out indirect effects through a mediator under randomization of treatment alone. We showed that IOTs can rule out indirect effects if the effect of treatment on the mediator can only move units in one direction, which is a very strong monotonicity assumption. Furthermore, IOTs cannot establish an indirect effect even if we assume a positive average effect of the mediator on the outcome. We derive sharp nonparametric bounds for the indirect effect under these various assumptions and estimate these bounds in two empirical applications. We also introduced a machine learning framework for evaluating the crucial monotonicity assumption. Finally, we also discussed how researchers can assess causal mechanisms when the assumptions needed to make inferences about indirect effects are questionable.
We believe there remains much to study in the field of causal mechanisms. In particular, future research should attempt to unify different accounts of “mechanisms” in a single broad framework. Currently, mechanisms are sometimes used interchangeably with “indirect effects,” but other times, scholars use mechanisms to refer to a theoretical construct that can explain an effect without reference to any mediation at all. In epidemiology, there is a tradition of defining causal mechanisms in the “sufficient cause framework,” where authors seek to find sets of variable levels that are sufficient to produce a particular outcome (VanderWeele, 2009). Connecting the ideas of mediation and mechanism to the theoretical underpinnings of the social sciences would help scholars develop better empirical tests of their theories.
@@ Appendix A Overview of the linear programming approach
Appendix B Bounds under Limited Defiers
The MMR assumption may be too strong in many cases, so we now derive bounds under a weaker restriction on the mechanism defiers. In particular, we assume the proportion of defiers is bound
$\rho_{01}=\sum_{y_{1}=0}^{1}\sum_{y_{0}=0}^{1}\psi_{y_{1}y_{0}01}\leq\gamma.$ (11)
This restriction can easily be incorporated into the linear programming approach described above.
Proposition 3.
Suppose that Assumption 1 and restriction (11) holds. Then, we have
$\max\left\{\begin{array}[]{l}-ATM-2\gamma\\ -p_{01\cdot 1}-p_{00\cdot 1}\\ -p_{00\cdot 1}-p_{10\cdot 0}-p_{00\cdot 0}\\ -ATM-p_{00\cdot 1}-\gamma\\ -p_{01\cdot 1}-p_{11\cdot 0}-p_{01\cdot 0}\\ -p_{01\cdot 1}-\gamma\end{array}\right\}\leq\bar{\delta}(1)\leq\min\left\{\begin{array}[]{l}ATM+2\gamma\\ p_{11\cdot 1}+p_{10\cdot 1}\\ p_{11\cdot 1}+\gamma\\ p_{11\cdot 1}+p_{11\cdot 0}+p_{01\cdot 0}\\ ATM+(1-p_{11\cdot 1})\\ ATM+p_{10\cdot 1}+\gamma\end{array}\right\}.\$ (12)
We derive these bounds under full generality about the relative size of $\gamma$ to observed data distribution, but we can provide more interpretable bounds if we assume that the proportion compliers is low relative to the obsered strata probabilities. In particular, if $\gamma\leq\min_{y,m}p_{ym\cdot 1}$, then we have
$\max\{-ATM-2\gamma,-p_{01\cdot 1}-\gamma\}\leq\bar{\delta(1)}\leq\min\{ATM+2\gamma,p_{11\cdot 1}+\gamma\}.$ (13)
One takeaway from these bounds is that when the ATM is zero (and the defiers group is small), the ANIE will be bounded between $-2\gamma$ and $2\gamma$. Substantively, this means that when there is no average effect of the treatment on the mediator, we can rule out indirect effects that are more twice the hypothesized size of the defier group. In other words, if the defier group is allowed to be half the size of the ATE, then a zero ATM would still be consistent with the possibility of an ANIE that explains the entire treatment effect.
Appendix C ANIE Bounds under Full Monotonicity
In this section, we formalize the full monotonicity assumption and derive the corresponding bounds on ANIE. We start by stating the assumption as follows:
Assumption 4 (Full Monotonicity).
$\mathbb{P}[M_{i}(1)\geq M_{i}(0)]=1,\mathbb{P}[Y_{i}(1,m)\geq Y_{i}(0,m)]=1,\mathbb{P}[Y_{i}(a,1)\geq Y_{i}(a,0)]=1$
Notice that full monotonicity assumption is strictly stronger than the MMR assumption. Moreover, we derive the ANIE bounds using the same linear programming procedure as described above:
Proposition 4.
Under Assumption 4, the sharp bounds for the ANIE are [^8] With straightforward algebraic manipulations, these bounds can be shown to be equivalent to those derived by Cai et al. (2008) using Balke’s algorithm.
$\begin{array}[]{l}\max\left\{\begin{array}[]{c}0\\ p_{10\cdot 1}-p_{10\cdot 0}\end{array}\right\}\leqslant\delta(0)\leqslant p_{10\cdot 1}+p_{11\cdot 1}-p_{10\cdot 0}\\ \max\left\{\begin{array}[]{c}0\\ p_{01\cdot 0}-p_{01\cdot 1}\end{array}\right\}\leqslant\delta(1)\leqslant p_{00\cdot 0}+p_{01\cdot 0}-p_{01\cdot 1},\end{array}$
This result shows that, under the full monotonicity assumption, the bounds on ANIE no longer cross zero.
Appendix D Additional MMR Falsification Tests
Figure SM.5: Validating the MMR Assumption (3) for the effect of museum visit on democratic supports (Balcells, Palanza and Voytas, 2022). The figure displays a histogram for estimated CATE values for the effect of $A_{i}$ on $M_{i}$ for each of the negative emotion mediators. CATEs are estimated using causal forests (Wager and Athey, 2018) with 2000 trees.